The initial temperature of the mixture was +20 °C. The mixture froze at –1.5 °C.

A total of 165 kJ of internal energy was transferred from the mixture to cool and freeze it.

specific heat capacity of the mixture = 3500 J/kg °C specific latent heat of fusion of the mixture = 255 000 J/kg

Calculate the mass of the mixture.

Give your answer to 2 significant figures.


1
Expert's answer
2021-04-13T06:36:03-0400

Express the amount of heat required for the process:


"Q=cm\Delta T+mL,\\\space\\
m=\frac{Q}{c\Delta T+L}=\frac{165000}{3500(-1.5-20)+225000}=1.1\text{ kg}."